# Explainable Malware Analysis: Concepts, Approaches and Challenges

### v1

@@ III-D Model-Specific Techniques for Post-Hoc Explainability
TreeSHAP and DeepSHAP represent two specialized implementations of SHAP grounded in the principles of Shapley values. TreeSHAP is tailored for tree-based models, offering a more efficient computation of exact SHAP values by operating in polynomial time, in contrast to the exponential time typically required by the general SHAP approach. In an illustrative application, Athanasiou et al. [14] leveraged TreeSHAP within an explainable risk prediction model for cardiovascular disease, utilizing this technique to furnish personalized explanations of the machine learning model’s predictions.
@@
Saliency Maps
They serve as critical tools in attribution analysis by showing the pixels that significantly influence image classification decisions. These gradient-based methods, designed specifically for neural network models, can facilitate an understanding of the features most relevant to a model’s output. Among such methods, Layer-wise Relevance Propagation (LRP) and DeepLIFT stand out by providing a framework to assign importance scores to different elements of a network, offering a detailed explanation of a model’s decision-making process. Specifically, LRP identifies the contribution of various parts of the input data towards the final decision, a technique effectively employed by Wang et al. [117] for dynamic and explainable malware detection. By pinpointing malicious code snippets, their approach enhances the interpretability of malware classifiers.
@@ IV Explainable Malware Classification and Detection Approaches
The field of malware analysis and detection through the lens of explainable machine learning has not been sufficiently explored in academic and practical research. The preceding section provided an introduction to the concept of explainability in ML, detailed various models and techniques for enhancing explainability, and reviewed relevant research in the broader field of explainable ML. Moving forward, this section will focus specifically on the application of explainable ML in the context of malware classification and detection. It will cover the diverse approaches researchers have developed for detecting and classifying malware, with a special emphasis on explainability. These approaches are organized by the types of target systems, including Windows Portable Executable (PE) files, hardware systems, Android devices, PDF documents, and Linux files. This classification is depicted in Figure 5, which provides a clear framework for understanding how explainable ML techniques are applied across different computing platforms to address malware threats.
Figure 5: Explainable Malware Classification and Detection Approaches [125].
@@ IV-A Windows PE-Based Malware Approaches
Windows PE is a file format based on the Common Object File Format (COFF) specification and holds significant importance within the Windows operating system family. As illustrated in Figure 6, the structure of a PE file starts with a header that was initially used by the MS-DOS operating system. When the executable is loaded, MS-DOS runs a stub program to ensure backward compatibility. Next, the COFF header provides detailed specifications of the executable file. It is followed by an optional header, which adds flexibility and supports future enhancements to the file structure. Following this, the section header divides the executable into distinct sections. These sections are made up of blocks of memory and support page swapping to address memory limitations, which leads to organizing the executable into structured segments for efficient execution. The Windows operating system has emerged as the predominant platform on personal computers, which has increased its vulnerability to malware attacks. Despite this risk, there has been limited research focused on explainable malware detection methodologies specifically for Windows.
Figure 6: Structure of Windows PE file.
Developing ML-based models that learn discriminative features from raw inputs requires a rigorous process of feature extraction, which is both time-consuming and complex. To address this challenge, Raff et al. [92] introduce ”MalConv,” a novel architecture for malware detection. This architecture leverages the entire executable as input for a CNN, which could distinguish this research from the limitations discussed in other studies surveyed. The MalConv architecture, illustrated in Figure 7, utilizes a methodology similar to those employed in speech and signal processing [43], text understanding [128], and image classification [111], where CNNs effectively extract pertinent features.
MalConv is designed as a static architecture for the identification of static malware, which combines CNN activation functions with global max-pooling before progressing to fully connected layers. This approach guarantees that the ML model generates activations independent of the features’ spatial locations, which leads to enhancing its ability to classify and detect malware. The discussion surrounding Windows PE malware detection includes techniques that are gradient-based, model-agnostic, and reliant on image representations. The forthcoming section will discuss an analysis of existing strategies for malware classification and detection, with Table I specifically addressing the application of model-agnostic, gradient-based, and image-based methodologies in the context of Windows PE-based malware research.
@@
Bose et al. 2020  [19] Analyzed the MalConv architecture and introduced a framework for analyzing malware detection networks end-to-end Their contribution involves gradient analysis representing the gradient embeddings of malicious files and interpolating between correctly classified samples to form a decision boundary between two classes to address the network’s decision in a better way Requires better interpretation of Neural Network results Gradient Analysis
Mathews 2019  [77] Identify potential vulnerabilities and flaws in feature engineering, which highlights the global characteristics in order to differentiate between two malware classes Introduced the explainability framework to classify the malware on windows PC Computed set of content-based features Proposed extraction of statistical features which reflects PE files Used LIME framework to explain window PC malware classification Required analysis and comparison of different model-agnostic explainable approaches LIME
Pirch et al. 2021  [90] Aims to provide an explainable CNN model that will predict tags for malware to be used for grouping and organizing the malware samples Each feature, also called a ’token,’ is assigned a relevance score which is used to determine how each of the tokens affected the tag that the model predicted Used two measurements to determine the quality of the explanation, descriptive accuracy, and descriptive sparsity Limited to behavioral-based approach LRP
Marais et al. 2021  [76] Their main focus is to interpret the models to reduce the False Positive (FP) rates Proposed malware detection model and their results are compared with LGBM, XGBoost, and DNN Employed the GradCam++ explainability approach on the CNN model to identify the important pixel that has the most influence on the CNN model’s prediction. Malware experts can use this tool to determine the nature of a suspicious file Lacks the attention mechanisms Limited to static-based approach, computational time need to be considered GradCam++
@@
Li et al. 2021  [63] Aims to provide an interpretable feed-forward neural network o interpret non-linear models They developed IFFNN with high accuracy and interpretability to detect malware Used MNIST dataset for qualitative interpretation of IFFNN The interpretation’s fidelity needs further evaluation and validation IFFNN
Chen et al. 2019  [24] Validate fidelity of explanations for dynamic malware analysis Validated the fidelity of interpretation by extending LIME to dynamic image-based malware classification. Developed two case studies to understand the advantages of vision-based interpretation frameworks. Missing explanation on the effectiveness of LIME for interpretation due to the fact that there are already better techniques LIME
Lin and Chang 2021  [67] Aims to implement an interpretable ensemble model for detecting image-based malware They designed a deep ensembled detector and deep taylor decomposition approach and interpreted the prediction results of the detector using LIME, SHAP and LRP explainability techniques. Required evaluation of more base classifiers to improve the detection accuracy Datasets containing the original malware files are not used to locate the semantic features LIME, SHAP, and LRP
Alani et al. 2023  [5] The paper focuses on detecting obfuscated malware on Windows systems using features extracted from memory dumps Introduces an ML-based system for detecting obfuscated Windows malware that achieves a detection accuracy of over 99% using only five features extracted from memory dumps. Provides detailed model explainability using SHAP, which enhances the transparency and trustworthiness of the system. The model relies on a limited set of features extracted from memory dumps. If malware evolves to avoid or alter these specific features, the model’s effectiveness may decrease. The high accuracy reported is based on the specific dataset used in the experiments (MalMem2022). The performance might vary with different datasets. SHAP
TABLE II: Research Addressing Explainable Machine Learning in Windows Malware (Continued).
Paper/Year Focus/Objective Contribution Limitation XAI technique
Ciaramella et al. 2024  [27] The paper proposes a DL-based model for detecting ransomware The paper introduces a DL approach that converts executable files into images and applies CNNs for ransomware detection. The model incorporates Grad-CAM to highlight the regions of the input images responsible for the model’s predictions, thereby improving the interpretability and explainability of the detection results. The focus is primarily on ransomware detection, and the model’s applicability to other forms of malware is not discussed. Grad-CAM
Anthony et al. 2024  [9] Develop an interpretable and robust ML model for malware detection using LENs The paper extends the application of LENs to the domain of malware detection using the large-scale EMBER dataset. It introduces a customized version of LENs that enhances the fidelity of logic explanations, which makes them more aligned with the model’s predictions. The process of generating FOL rules and optimizing these explanations requires considerable computational resources. This high computational cost can limit the scalability of LENs, making them less practical for real-time application. LENs
Gulmez et al. 2024  [44] This study aims to develop an explainable DL-based system named XRan that detects ransomware using dynamic analysis. XRan enriches the feature space by integrating API call sequences, DLL sequences, and Mutex sequences. Utilizes XAI to provide local and global explanations, improving the interpretability of the model. The dynamic analysis phase of XRan is resource-intensive and time-consuming compared to static analysis. This limitation is inherent to all dynamic analysis-based techniques, but it particularly affects XRan due to its comprehensive feature extraction process that includes API calls, DLL sequences, and Mutex sequences. LIME, and SHAP
Aryal et al. 2024  [12] It aims to improve the success of adversarial evasion attacks on malware detectors by using XAI techniques to identify and target critical regions of Windows PE malware files for perturbation. Utilizes SHAP values to identify optimal regions in Windows PE malware for effective adversarial perturbation. Demonstrates improved evasion rates of malware detectors using SHAP-guided perturbations. This paper focuses on the effectiveness of SHAP-guided adversarial perturbations without fully exploring how these perturbations might perform across different types of Windows PE malware, such as trojans, ransomware, worms, and others. SHAP
Ghadekar et al. 2024  [12] It aims to develop an advanced DL approach for multi-class malware detection with the goal of enhancing feature extraction capabilities and integrating XAI principles. The paper introduces a modified GNN (deeperGCN) that significantly improves malware detection accuracy. It Integrates XAI using GradCAM to provide transparent and interpretable insights into the model’s decision-making process. The process of converting byte and ASM files into images requires significant preprocessing steps, including feature extraction and transformation into graphical representations. This preprocessing is complex and time-consuming, potentially limiting the model’s applicability in environments where rapid detection is important. GradCAM
@@ IV-A1 Gradient-Based Approach
The gradient-based methodology measures the impact of input features on predictions by assigning weights to different parts of an executable. To elucidate the decision-making processes of Deep Neural Networks (DNNs), Bose et al. [19] examine the MalConv architecture using the open-source ’emberMalConv’ framework. This study seeks to understand how the architecture distinguishes between malicious and benign executables based on their raw data. Ember, which is a tool utilized for training static PE malware models within the ML domain, highlights the MalConv architecture’s unique ability to attribute significant weight to specific executable parts, thereby having a significant influence on the classification results. Their research introduces a sophisticated framework based on gradient analysis, which maps gradient embeddings from malicious files and interpolates between accurately classified instances to define a clear decision boundary between categories. By analyzing the interpolation among samples, the study explores filter activations to investigate if there is a connection between different filter pairs. This leads to the development of a correlation heatmap for the filters, providing insights into how they interact. Remarkably, one filter specializes in identifying malicious traits within a file, and another filter focuses on generalizing these findings across various samples. The proposed framework transcends the MalConv model, offering a general method suitable for classification tasks in any neural network.
@@ IV-A2 Model-Agnostic-Based Approach
The Model-Agnostic approach employs post-hoc explainability techniques that can be applied across any opaque or black-box model. This method clarifies the predictions by simplifying the complex original model into a more understandable local surrogate model. The research presented by Mathews [77] introduces an explainability framework aimed at classifying two distinct malware families on Windows PCs. Firstly, the researchers calculate content-based features and extract statistical features derived from Hex and assembly views. These features are indicative of the PE file’s structure.
Their investigation shows shortcomings in the feature selection process and emphasizes the global characteristics through which a model learns to distinguish between the two malware categories. To elucidate the outcomes produced by the deep learning model, they utilize the LIME explainability framework.
@@
Figure 8 presents an integrated model for dynamic malware detection based on the principles of explainable ML, combining dynamic analysis with the training of a surrogate learning model using malware tags. This model leverages explainable ML techniques to clarify the surrogate model’s predictions and behavioral patterns, aiming to enhance understanding of its decision-making process. Another research [90] seeks to develop a CNN model that accurately predicts malware tags, which are essential for categorizing and organizing malware samples and makes the prediction process transparent. This endeavor involves a thorough dynamic analysis that examines malware tags to inform the training of the surrogate learning model. The study progresses by applying explainable machine learning to demystify how the CNN model arrives at its predictions. Each detected feature, referred to as a ’token’ in the research, receives a relevance score that indicates its impact on the predicted malware tag. To assess the quality of these explanations, the authors employ two measures: descriptive accuracy, which evaluates the precision with which an explanation captures the influential features of a prediction, and descriptive sparsity, which identifies the superfluous features within these explanations. The model’s effectiveness is validated through its performance in classifying three types of tags—sandbox, family, and clustering—with each category achieving an accuracy rate of over 90%.
Figure 7: MalConv architecture [92].
@@
In a study, Alani et al. [5] present an ML-based system designed to detect obfuscated malware on Windows platforms with high accuracy and efficiency. The system utilizes a variety of classifiers, including random forest (RF), logistic regression (LR), decision tree (DT), Gaussian Naive Bayes (GNB), and extreme gradient boosting (XGB). Through an evaluation process, XGB was identified as the best-performing classifier. Moreover, The system relies on features extracted from memory dumps using the VolMemLyzer tool. The feature selection algorithm, i.e., Recursive Feature Elimination (RFE), identifies the five most effective features, resulting in a streamlined model that maintains an accuracy rate exceeding 99%. The selected features include the total number of services, average number of dynamic-link libraries (DLLs) per process, total number of mutant handles, number of kernel drivers, and shared process services. The system’s detection capabilities are bolstered by its explainability, achieved through SHAP. SHAP values provide insight into the impact of each feature on the model’s predictions. The evaluation of the system demonstrates its high accuracy and rapid detection speed, with a processing time of 0.413 microseconds per instance. Despite its robustness, the paper has some potential limitations, such as the model’s dependence on specific features and vulnerability to adversarial attacks.
Figure 8: Explainable malware method for Tagvetting.

### v2

@@ III-D Model-Specific Techniques for Post-Hoc Explainability
TreeSHAP and DeepSHAP. They represent two specialized implementations of SHAP grounded in the principles of Shapley values. TreeSHAP is tailored for tree-based models, offering a more efficient computation of exact SHAP values by operating in polynomial time, in contrast to the exponential time typically required by the general SHAP approach. In an illustrative application, Athanasiou et al. [62] leveraged TreeSHAP within an explainable risk prediction model for cardiovascular disease, utilizing this technique to furnish personalized explanations of the machine learning model’s predictions.
@@
Saliency Maps. They serve as critical tools in attribution analysis by showing the pixels that significantly influence image classification decisions. These gradient-based methods, designed specifically for neural network models, can facilitate an understanding of the features most relevant to a model’s output. Among such methods, Layer-wise Relevance Propagation (LRP) and DeepLIFT stand out by providing a framework to assign importance scores to different elements of a network, offering a detailed explanation of a model’s decision-making process. Specifically, LRP identifies the contribution of various parts of the input data towards the final decision, a technique effectively employed by Wang et al. [66] for dynamic and explainable malware detection. By pinpointing malicious code snippets, their approach enhances the interpretability of malware classifiers.
@@ IV Explainable Malware Classification and Detection Approaches
The preceding section provided an introduction to the concept of explainability in ML, detailed various models and techniques for enhancing explainability, and reviewed relevant research in the broader field of explainable ML. Moving forward, this section will focus on the application of explainable ML in the context of malware classification and detection. These approaches are organized by the types of target systems, including Windows PE files, hardware systems, Android devices, PDF documents, and Linux files. This classification is depicted in Fig 4, which provides a clear framework for understanding how explainable ML techniques are applied across different computing platforms to address malware threats.
By categorizing existing explainable malware detection strategies by the target platform—Windows PE, Android, hardware, PDF, and Linux— we highlight how explainability methods adapt to each system’s unique execution environment and feature set. We also provide comprehensive tables (II, I, III, IV) summarizing key contributions, limitations, and XAI techniques. These tables are provided for quick reference. However, we will make explanatory context in each subsection.
Figure 4: Explainable Malware Classification and Detection Approaches
@@ IV-A Windows PE-Based Malware Approaches
Windows is the most widely used desktop OS, making Windows PE files a significant target for malware. Below, we summarize the PE format’s structure and then highlight state-of-the-art studies based on key explainable detection methods grouped by the type of XAI approach (gradient-based, model-agnostic, and image-based). Finally, Table I lists major works in Windows PE-based malware detection, comparing their focus, contributions, limitations, and XAI techniques.
Windows PE is a file format based on the Common Object File Format (COFF) specification and holds significant importance within the Windows operating system family. The structure of a PE file starts with a header initially used by the MS-DOS operating system. When the executable is loaded, MS-DOS runs a stub program to ensure backward compatibility. Next, the COFF header provides detailed specifications of the executable file. It is followed by an optional header, which adds flexibility and supports future enhancements to the file structure. Following this, the section header divides the executable into distinct sections. These sections comprise blocks of memory and support page swapping to address memory limitations, which leads to organizing the executable into structured segments for efficient execution.
The Windows operating system has emerged as the predominant platform on personal computers, which has increased its vulnerability to malware attacks. Despite this risk, limited research has focused on explainable malware detection methodologies specifically for Windows.
Developing ML-based models that learn discriminative features from raw inputs requires feature extraction, which is time-consuming and complex. To address this challenge, Raff et al. [68] introduce ”MalConv,” a novel architecture for malware detection. This architecture leverages the entire executable as input for a CNN. The MalConv architecture utilizes a methodology similar to techniques employed in speech and signal processing [69], text understanding [70], and image classification [71], where CNNs effectively extract pertinent features.
MalConv is designed as a static architecture for identifying static malware, which combines CNN activation functions with global max-pooling before progressing to fully connected layers. This approach guarantees that the ML model generates activations independent of the features’ spatial locations, which leads to enhancing its ability to classify and detect malware. The discussion surrounding Windows PE malware detection includes gradient-based, model-agnostic techniques, and reliance on image representations.
The forthcoming section will discuss an analysis of existing strategies for malware classification and detection, with Table I explicitly addressing the application of model-agnostic, gradient-based, and image-based methodologies in the context of Windows PE-based malware research.
Table I summarizes research on explainable Windows malware detection. Focus/Objective clarifies the primary goal, the Contribution outlines each paper’s novelty, and the XAI Technique highlights the interpretability approach used.
@@
Bose et al. 2020 [72] Analyzed MalConv and proposed an end-to-end gradient-based explanation framework Introduced gradient analysis mapping malicious file embeddings and interpolating between correctly classified samples to clarify decision boundaries Limited interpretability of Neural Network decisions Gradient Analysis
Mathews 2019 [73] Identify feature-engineering flaws and global traits distinguishing malware classes Proposed an explainability framework using content-based and PE-file statistical features, with LIME for model interpretation Lacks comparison with other model-agnostic XAI methods LIME
Pirch et al. 2021 [74] Develop an explainable CNN to predict malware tags for clustering and organization Assigns a relevance score to each feature (“token”) for tagging malware, evaluated by descriptive accuracy and sparsity Restricted to a behavioral-based approach LRP
Marais et al. 2021 [75] Improve model interpretability to reduce false positives Developed a CNN-based malware model compared with LGBM, XGBoost, and DNN, using GradCam++ to highlight critical pixels for reduced false positives Missing attention mechanisms; limited to static analysis and higher computation time GradCam++
@@
Li et al. 2021 [76] Introduce an interpretable feed-forward neural network (IFFNN) for malware detection Proposed IFFNN with high accuracy and interpretability; tested with MNIST for qualitative assessment Interpretation fidelity requires more validation IFFNN
Chen et al. 2019 [77] Validate explanation fidelity in dynamic malware analysis Extended LIME for vision-based interpretation in dynamic malware analysis, with two case studies showing interpretability benefits Does not compare LIME with more advanced XAI methods LIME
Lin and Chang 2021 [78] Implement a deep ensemble detector for image-based malware with interpretable outputs Proposed Selective Deep Ensemble Learning (SDEL) and used LIME, SHAP, and LRP to interpret image-based malware predictions It needs more base classifiers and original malware files to locate semantic features LIME, SHAP, and LRP
Alani et al. 2023 [79] Detect obfuscated Windows malware from memory dumps using minimal features Achieves ¿99% detection using only five memory-dump features, with SHAP-based interpretability Dependent on a limited feature set and dataset-specific performance SHAP
Ciaramella et al. 2024 [80] Proposes a CNN-based approach for ransomware detection using image conversion Uses Grad-CAM to visualize high-impact image regions, improving ransomware detection interpretability Focused mainly on ransomware, with unclear generalization to other malware Grad-CAM
Anthony et al. 2024 [81] Adapt LENs for interpretable large-scale malware detection Provides high-fidelity FOL rule explanations, balancing accuracy with transparency Generating/optimizing FOL rules is computationally heavy, limiting real-time application LENs
Gulmez et al. 2024 [82] Propose XRan, an XAI-enabled CNN approach combining API, DLL, and Mutex features for ransomware detection Combines API/DLL/Mutex features in a two-layer CNN, reaching 99.4% TPR with LIME/SHAP explanations High overhead for dynamic feature extraction LIME, and SHAP
Aryal et al. 2024 [83] Use SHAP to locate and perturb critical regions of Windows PE malware for enhanced evasion Demonstrates SHAP-guided perturbation boosts adversarial evasion rates while preserving malware functionality Limited to specific Windows PE samples; broader malware families not tested SHAP
Ghadekar et al. 2024 [84] Propose a deeperGCN model for multi-class malware detection with integrated GradCAM explainability Reaches 97% accuracy on combined byte/ASM images; GradCAM highlights key regions Preprocessing (byte/ASM to image) is time-consuming and may limit real-time use GradCAM
@@ IV-A1 Gradient-Based Approach
The gradient-based methodology measures the impact of input features on predictions by assigning weights to different parts of an executable. To elucidate the decision-making processes of Deep Neural Networks (DNNs), Bose et al. [72] examine the MalConv architecture using the open-source ’emberMalConv’ framework. This study seeks to understand how the architecture distinguishes between malicious and benign executables based on their raw data. Ember, which is a tool utilized for training static PE malware models within the ML-based domain, highlights the MalConv architecture’s ability to attribute significant weight to specific executable parts, thereby having a significant influence on the classification results.
Their research introduces a sophisticated framework based on gradient analysis, which maps gradient embeddings from malicious files and interpolates between accurately classified instances to define a clear decision boundary between categories. By analyzing the interpolation among samples, the study explores filter activations to investigate if there is a connection between different filter pairs. This leads to the development of a correlation heatmap for the filters, providing insights into how they interact. One filter specializes in identifying malicious traits within a file, and another filter focuses on generalizing these findings across various samples. The proposed framework transcends the MalConv model and offers a general method suitable for classification tasks in any neural network.
In summary, gradient-based methods for Windows PE malware detection effectively pinpoint which bytes or segments of an executable are most influential for classification, offering highly granular explanations. However, they often require large labeled datasets and can be susceptible to adversarial manipulation if attackers target the most salient bytes. Despite these limitations, gradient-based XAI remains a powerful tool when fine-grained feature importance is crucial.
@@ IV-A2 Model-Agnostic-Based Approach
This method clarifies the predictions by simplifying the complex original model into a more understandable local surrogate model. The research presented by Mathews [73] introduces an explainability framework aimed at classifying two distinct malware families on Windows PCs. Firstly, they calculate content-based features and extract statistical features derived from Hex and assembly views. These features are indicative of the PE file’s structure. Their investigation shows shortcomings in the feature selection process and emphasizes the global characteristics through which a model learns to distinguish between the two malware categories. To elucidate the outcomes produced by the deep learning model, they utilize the LIME framework.
@@
A study by Pirtch et al. [74] seeks to develop a CNN model that accurately predicts malware tags. This work involves a thorough dynamic analysis that examines malware tags to inform the training of the surrogate learning model. Each detected feature, referred to as a ’token’ in the research, receives a relevance score that indicates its impact on the predicted malware tag. To assess the quality of these explanations, the authors employ two measures: descriptive accuracy, which evaluates the precision with which an explanation captures the influential features of a prediction, and descriptive sparsity, which identifies the superfluous features within these explanations. The model’s effectiveness is validated through its performance in classifying three types of tags—sandbox, family, and clustering—with each category achieving an accuracy rate of over 90%.
@@
In a study, Alani et al. [79] present an ML-based system designed to detect obfuscated malware on Windows platforms with high accuracy and efficiency. The system utilizes a variety of classifiers, including RF, logistic regression, decision trees, Gaussian Naive Bayes (GNB), and extreme gradient boosting (XGB). Through an evaluation process, XGB was identified as the best-performing classifier. Moreover, The system relies on features extracted from memory dumps using the VolMemLyzer tool. The feature selection algorithm, i.e., Recursive Feature Elimination (RFE), identifies the five most effective features, resulting in a streamlined model that maintains an accuracy rate exceeding 99%. The selected features include the total number of services, average number of dynamic-link libraries (DLLs) per process, total number of mutant handles, number of kernel drivers, and shared process services. The system’s detection capabilities are bolstered by its explainability, achieved through SHAP. SHAP values provide insight into the impact of each feature on the model’s predictions. The evaluation of the system demonstrates its high accuracy and rapid detection speed, with a processing time of 0.413 microseconds per instance. Despite its robustness, the paper has some potential limitations, such as the model’s dependence on specific features and vulnerability to adversarial attacks.
@@
Overall, model-agnostic frameworks (e.g., LIME, SHAP, LENs) are popular for Windows PE malware detection thanks to their flexibility: they can explain virtually any classifier. Yet, these post-hoc explanations can vary from sample to sample, sometimes lacking global consistency. Their simplicity and model independence, however, make them valuable for real-world malware detection pipelines.